Hello!
if (x-2)squared = 25 and x < 0, what is the value of x?

Found 2 solutions by ewatrrr, lwsshak3:
Answer by ewatrrr(24785)

(x-2)^2 = 25 and x < 0
(x-2) = -5
x = -3

Answer by lwsshak3(11628)

if (x-2)squared = 25 and x < 0, what is the value of x?
(x-2)^2=25
x-2=±√25
x=2±5
x=2-5
x=-3